OpenStudy (anonymous):

4X - 5Y = -18 , 5X + 4Y = -2

OpenStudy (anonymous):

Solve for x in terms of y in one equation, then substitute that expression for x in second equation and solve for y. Plug the solution for y back into you expression for x and solve for x. Check that your solution satisfies each equation!

OpenStudy (anonymous):

4X - 5Y = -18.............1 multiply by 5 5X + 4Y = -2...............2 multiply by 4 20x-25y =90...................1 20x+16 y = -8...................2 multiply by -1 20x-25y =90...................1 -20x-16y=8.......................2 add equation 1 to equation 2 -41y=98 y=-2.39 sub value of y in any equation 4x-5(-2.39)=-18 x=-7.4875 you are welcome .

OpenStudy (anonymous):

waleed made an error: 5 x equation 1 should be = -90, not 90. Sign was wrong. Good luck.

OpenStudy (anonymous):

yeah .. you are right ! .... thnx ... I do mistake all time .. one day i will be better 4X - 5Y = -18.............1 multiply by 5 5X + 4Y = -2...............2 multiply by 4 20x-25y = - 90...................1 20x+16 y = -8...................2 multiply by -1 20x-25y = - 90...................1 -20x-16y = 8.......................2 add equation 1 to equation 2 -41y=-82 y = 2 sub value of y in any equation 4x-5(2)=-18 x=-2